# Common Corpus: The Largest Collection of Ethical Data for LLM Pre-Training

### v1

@@ B.2 Open Culture
Table 6: Subsets of Open Culture with language coverage, domains, and token counts.
@@
  • German PD. This dataset contains texts from various sources, including the Mannheim Corpus of Historical Newspapers and Magazines [^55] https://repos.ids-mannheim.de/fedora/objects/clarin-ids:mkhz1.00000/datastreams/CMDI/content (Mannheimer Korpus Historischer Zeitungen und Zeitschriften). This dataset is made up of 21 German newspapers and magazines. The texts were originally published between 1737 and 1905. The corpus was originally digitized between 2009 and 2011. The corpus was made available by the Institut für Deutsche Sprache in 2013.
@@
B.3 Open Science
In Table 7, we present the total token counts per collection inside of the Open Science part of Common Corpus.
@@ B.4 Open Code
Appendix C Open Culture Verification
@@
Appendix D Cleaning and Curation
D.1 Text Segmentation
@@
In this respect, the insurance business investment portfolio can be considered conservatively managed as it is largely composed of corporate, sovereign, and supranational bonds, term loans as well as demand deposits. Following the previous year, the group continued to diversify its holdings into investment-grade corporate bonds. It should be noted that bonds and term loans are held to maturity in accordance with the group’s business model policy of "inflows".
@@
The income and expenses recognized for the insurance contracts issued by the group appear in the income statement in "Net income of other activities" and "Net expense of other activities".
@@
The group adopts a "prudent approach" to its management of the risks to which it could be exposed through its insurance activities. Risk of counterparty. As stated above, insurance companies only invest in assets (bank deposits, sovereign bonds, supranational agencies, or corporate bonds).
@@
D.2 OCR Error Detection
OCRoscope.
To illustrate this approach, this long text is correctly identified as French with >99% confidence by cld2, as despite the many mistakes, there are enough non-ambiguous French words:
NOUVELLES POLI TI QÛ E S. Suede. Stockholm , le 2 5 décembre 1792. Le général Toll ira à Varsovie en quarté d’envoyé de la Suede auprès du roi et de la république ; A 1 même rey.u l’ordre de s’y rendra incessamment. 11 paraît que k Uc-régeik a des craintes ; il a fait venir chez lji les membres c Ij“‘ tribunal 4e la cour , et leur a rtmis son lesfca n at. La fermentation qu’a causée 1 ,’ari r?tavh n k M p v riote Thorild tî’est pas appaisée y le luigage qv’il a yailé an duc-régent a été bien entendu par le peu) k y ir M» (U i n’entendrait pas l’apostrdphe suivante ? ttRxc3xa7nd >la libuk à r otre raison , et ne et nous force pas de i’ache’ef r i te n :e sang,.
@@ D.2 OCR Error Detection
Yet one short n-gram ("n k M p v riote Thorild") is classified as unknown by cld2.
OCRerrcr.
The following is a low-error example sentence taken from Common Corpus:
@@
And the OCRerrcr detection (with formatting for clarity):
They did not approach <<<er>>>cer,<<</er>>> but turned away and passed <<<er>>>irom<<</er>>> her presence, filled with sorrow and moved with sympathy, which her intense emotions seemed to communicate to even these thoughtless young men of the <<<er>>>tho<<</er>>> plains.
D.3 OCR Correction
Here is an example of text containing various OCR errors:
@@ D.3 OCR Correction
And here is the text corrected by our model, OCRonos:

### v2

@@ D.2 Open Culture
Table 8: Subsets of Open Culture with language coverage, type of document, and token count.
@@
  • German PD. This dataset contains texts from various sources, including the Mannheim Corpus of Historical Newspapers and Magazines [^35] https://repos.ids-mannheim.de/fedora/objects/clarin-ids:mkhz1.00000/datastreams/CMDI/content (Mannheimer Korpus Historischer Zeitungen und Zeitschriften). This dataset is made up of 21 German newspapers and magazines. The texts were originally published between 1737 and 1905. The corpus was originally digitized between 2009 and 2011 and made available by the Institut für Deutsche Sprache in 2013.
@@
D.3 Open Science
In Table 9, we present the token counts inside of the Open Science collections.
@@ D.4 Open Code
Appendix E Open Culture Verification
@@
Appendix F Cleaning and Curation
We release the cleaning and curation tools as part of our Bad Data Toolbox, a suite designed to make challenging digitized resources usable for downstream NLP tasks and large-scale corpus construction.
F.1 Text Segmentation: Segmentext
Repository: [Uncaptioned image] PleIAs/Segmentext
Segmentext is a token classification model trained for the structural segmentation of documents, with a specific focus on robustness to digitisation artefacts and broken or unstructured input text. In contrast to layout-based segmentation approaches that operate on visual signals such as bounding boxes or font information, Segmentext works exclusively on raw character sequences, reconstructing editorial structure from text alone. This makes it applicable in post-OCR pipelines where layout information has been lost.
The model is a DeBERTa-v2 (He et al., 2021) variant for token classification, fine-tuned on 3,500 manually annotated documents. The training data was drawn from three datasets spanning complementary domains: Open Government, Open Culture, and Open Science. This domain diversity is intended to give the model broad applicability across document types and languages.
Segmentext supports the following text segmentation:
  • Text: the text itself.
  • Title: any type of heading.
  • Table: tabular content.
  • Separator: a segmentation separator, mostly based on newline, with some variations due to text segmentation understanding.
  • Dialog: any kind of speaker-attributed intervention.
  • Bibliography: statement of a specific bibliographic reference, either in a bibliography section or a footnote.
  • Contact: personal information, which can be especially useful in the context of PII removal.
  • Paratext: any non-meaningful text included in standard documents like headers, page numbering, section recall, etc.
  • Author: author names and signatures.
  • Date: statement of date and time, common in letters and newspaper articles.
  • Keyword: list of keywords, especially common in scientific publications.
@@ F.1 Text Segmentation: Segmentext
In this respect, the insurance business investment portfolio can be considered conservatively managed as it is largely composed of corporate, sovereign, and supranational bonds, term loans as well as demand deposits. Following the previous year, the group continued to diversify its holdings into investment-grade corporate bonds. It should be noted that bonds and term loans are held to maturity in accordance with the group’s business model policy of ”inflows”.
@@
The income and expenses recognized for the insurance contracts issued by the group appear in the income statement in ”Net income of other activities” and ”Net expense of other activities”.
@@
The group adopts a ”prudent approach” to its management of the risks to which it could be exposed through its insurance activities. Risk of counterparty. As stated above, insurance companies only invest in assets (bank deposits, sovereign bonds, supranational agencies, or corporate bonds).
@@
F.2 OCR Error Detection: OCRoscope
Repository: [Uncaptioned image] PleIAs/OCRoscope
OCRoscope is a lightweight Python library for estimating the quality of OCR-processed text without access to a ground-truth reference. Rather than relying on character- or word-level edit distance against a gold standard, OCRoscope exploits the differential behaviour of language detection models across text granularities. On long texts, language detection is robust to noise; on short character sequences, it degrades rapidly in the presence of OCR errors. The library leverages this property by applying language detection (via pycld2) to rolling 7-grams extracted from the input text and comparing the per-gram language assignments against the document-level language detected on the full text. The proportion of n-grams assigned to a mismatching or unknown language yields the OCR quality score: a value close to 1 indicates high-quality text, while lower values signal increasing levels of digitisation noise.
OCRoscope returns two primary metrics: a standardised OCR quality rate and a standardised rate of non-character content (punctuation, digits, whitespace anomalies). The tool supports over 80 languages via cld2’s language coverage. The choice of cld2 over more accurate successors such as cld3 is intentional: cld2’s greater sensitivity to short noisy sequences makes it a more discriminating signal for OCR quality estimation, yielding a score distribution that is well spread across the quality spectrum and therefore easier to threshold for downstream filtering decisions.
The tool is designed for pipeline use cases in which large corpora must be rapidly triaged: it provides a fast, reference-free quality signal that can guide decisions about whether texts warrant OCR correction, manual review, or outright exclusion.
Below we illustrate OCRoscope performance on this long text, which is correctly identified as French with >>99% confidence by cld2. Despite the many mistakes, there are enough non-ambiguous French words:
NOUVELLES POLI TI QˆU E S. Suede. Stockholm , le 2 5 décembre 1792. Le général Toll ira à Varsovie en quarté d’envoyé de la Suede auprès du roi et de la république ; A 1 même rey.u l’ordre de s’y rendra incessamment. 11 paraît que k Uc-régeik a des craintes ; il a fait venir chez lji les membres c Ij“‘ tribunal 4e la cour , et leur a rtmis son lesfca n at. La fermentation qu’a causée 1 ,’ari r?tavh n k M p v riote Thorild tî’est pas appaisée y le luigage qv’il a yailé an duc-régent a été bien entendu par le peu) k y ir M (U i n’entendrait pas l’apostrdphe suivante ? ttRxc3xa7nd >>la libuk à r otre raison , et ne et nous force pas de i’ache’ef r i te n :e sang,.
@@ F.2 OCR Error Detection: OCRoscope
Yet one short n-gram (“n k M p v riote Thorild”) is classified as unknown by cld2. Complete processing by OCRoscope yielded a 41% rate of non-recognized 7-grams, resulting in an OCR quality rate of 59%. In comparison, the self-estimated rate of OCR-validated words by the French National Library is significantly higher (85%) for the whole document.
F.3 OCR Error Detection: OCRerrcr
Repository: [Uncaptioned image] PleIAs/OCRerrcr
OCRerrcr is a small token classification model trained to identify individual OCR errors at the token level within a text. Unlike OCRoscope, which produces a document-level quality score, OCRerrcr generates a fine-grained annotation of which tokens are likely to be erroneous, making it suitable for targeted correction workflows and for producing spatially explicit error rate estimates.
The model is a DeBERTA-v3 (He et al., 2023) with 400M parameters fine-tuned on a corpus of approximately 1,000 documents with manually labelled OCR errors, drawn from Open Government and Open Culture. This dual-domain training is intended to give the model coverage across both modern administrative language and historical or literary language, which present different error profiles. OCRerrcr provides an accurate reference-free OCR error rate estimate, outperforming OCRoscope in precision, particularly on documents with a low or moderate error rate. However, it scales less efficiently than OCRoscope to very large corpora due to its per-token inference cost.
The model’s name is itself a playful illustration of the problem it addresses: OCRerrcr is a plausible OCR misreading of OCRerror, with the vowels dropped as they might be in a degraded scan.
The following is a low-error example sentence taken from Open Culture:
@@ F.3 OCR Error Detection: OCRerrcr
The version with the errors highlighted by OCRerrcr is presented below:
They did not approach <<er>>cer,<</er>> but turned away and passed <<er>>irom<</er>> her presence, filled with sorrow and moved with sympathy, which her intense emotions seemed to communicate to even these thoughtless young men of the <<er>>tho<</er>> plains.
F.4 OCR Error Correction: OCRonos
Repository: [Uncaptioned image] PleIAs/OCRonos
OCRonos is a generative language model trained for the correction of badly digitised text, including OCR errors, incorrect word splits and merges, and broader structural artefacts introduced during digitisation. It belongs to the corrective tier of the Bad Data Toolbox, operating downstream of quality detection tools such as OCRoscope and OCRerrcr.
The training data for OCRonos consists of a diverse multilingual set of OCR-processed texts drawn from Open Culture and Open Government. The model is fine-tuned from Llama-3-8B (Grattafiori et al., 2024).
OCRonos is designed to be conservative: it is generally faithful to the original source material and rarely rewrites words that are not erroneous. On highly degraded inputs, however, it can function as a synthetic rewriting tool rather than a strict corrector. A known issue at scale is the occasional inclusion of spuriously repeated words, which can be filtered in post-processing. The model also mitigates a common failure mode observed in smaller generalist LLMs, namely language switching: faced with heavily noisy input, models such as GPT-3.5 or Claude Haiku have been observed to drift into a different language or script during correction. OCRonos is specifically trained to resist this behaviour.
Below we present an exemplar text containing various OCR errors:
@@ F.4 OCR Error Correction: OCRonos
The following is the version corrected by OCRonos:
@@
F.5 Toxicity Detection: Celadon
Repository: [Uncaptioned image] PleIAs/celadon
Celadon is a multilingual toxicity classifier designed specifically for the challenges posed by public-domain pre-training data: historical language, OCR noise, and multilingual content spanning nine languages. It was developed as part of the Toxic Commons project, described in Arnett et al. (2024), and was applied to filter harmful content in Common Corpus prior to its release.
The model is a fine-tune of DeBERTa-v3-small (He et al., 2023) with five independent classification heads, each corresponding to one dimension of toxicity: race and origin-based bias (including racism, xenophobia, and bias against immigration status), gender and sexuality-based bias (including sexism, homophobia, and transphobia), religious bias, ability bias (bias relating to physical, mental, or intellectual disability), and violence and abuse (graphic violence, threats, and incitement). Each head is a separate linear layer with a custom weighted cross-entropy loss function to handle class imbalance.
The model was trained on 600,000 samples drawn from ToxicCommons, a dataset of approximately 2 million annotated public-domain text segments from 13 multilingual cultural heritage collections. Annotations were generated in two stages: a small set of human annotations was used to calibrate a prompt for the instruct version of Llama 3.1-8B (Grattafiori et al., 2024), which then produced annotations at scale across five toxicity dimensions using a 0–3 severity scale. The training set was balanced to equalize non-toxic and toxic samples.
Celadon’s key design rationale is efficiency: it achieves toxicity detection performance comparable to Llama 3.1-8B-Instruct while being over 40 times faster (5 minutes versus 3.4 hours per 100,000 samples), making it practical for large-scale corpus filtering. The model is optimised for the specific challenges of its target domain: historical text, OCR noise, and non-English content, and may not generalise without adaptation to web-text corpora, for which it was not designed.
Appendix G Evaluations
Figure 4 presents the distribution of six document-level quality and character composition metrics across the five open types in a 300,000-document sample of Common Corpus:
  • Character Repetition measures the proportion of character pairs belonging to repeated character bigram sequences within a document. This value is expected to be higher for code data, due to specific syntax and varied punctuation.
  • Type-Token Ratio is the ratio of unique word forms to the total number of tokens. It captures lexical diversity; documents with values close to 1 are typically very short, while low values can indicate highly repetitive text. For Open Government, this value is expected to be lower as regulatory texts usually contain fixed, distinct terms and repetitive formulations.
  • Top Word Proportion is the share of all tokens accounted for by the single most frequent word in a document. This is another measure of linguistic difersity.
  • Alphanumeric Characters denotes the fraction of characters that are letters or digits. This proportion might be lower in code due to reliance on punctuation and whitespace symbols.
  • Whitespace Characters gives the proportion of whitespace characters, which tends to be higher in source code and tabular data than in running prose.
  • Uppercase Characters measures the share of uppercase letters and separates Open Source content, which is densely populated with identifiers, constants, and type annotations, from Open Culture and Open Government documents that follow standard typographic conventions.
In Tables 111213, and 14, we present per-language scores for the studied benchmarks. On MultiBLIMP, most of the scores are significantly above random (0.5); therefore, we also highlight the best and second-best scores.
Model Ours Gemma 3 XGLM BLOOM Ours Gemma 3 XGLM OLMo
350M 270M 564M 560M 1.2B 1B 1.7B 1B
abk 0.550 0.750 0.475 0.525 0.675 0.675 0.325 0.800
aln 0.733 0.755 0.709 0.700 0.728 0.750 0.675 0.690
amh 0.946 0.929 0.911 0.973 1.000 0.955 0.848 0.964
apu 0.964 0.964 0.893 0.786 0.964 0.929 0.964 0.893
aqz 0.214 0.357 0.429 0.500 0.429 0.429 0.714 0.214
arb 0.877 0.913 0.895 0.923 0.900 0.951 0.887 0.782
azz 0.734 0.744 0.729 0.720 0.729 0.758 0.773 0.686
bel 0.799 0.795 0.574 0.608 0.853 0.896 0.577 0.611
ben 0.571 0.762 1.000 0.810 0.762 0.905 0.857 0.524
bho 0.676 0.647 0.588 0.588 0.706 0.794 0.618 0.588
bor 0.722 0.631 0.697 0.610 0.697 0.627 0.680 0.668
bre 0.942 0.815 0.554 0.604 0.938 0.946 0.615 0.685
bua 0.680 0.718 0.670 0.718 0.670 0.641 0.699 0.660
bul 0.872 0.880 0.969 0.623 0.897 0.945 0.976 0.735
cat 0.885 0.852 0.961 0.950 0.919 0.931 0.953 0.735
ces 0.824 0.808 0.579 0.597 0.858 0.891 0.603 0.668
chu 0.670 0.648 0.582 0.635 0.659 0.663 0.593 0.632
cym 0.771 0.730 0.633 0.611 0.828 0.796 0.610 0.796
dan 0.980 1.000 0.840 0.800 0.980 1.000 0.740 0.940
deu 0.967 0.949 0.961 0.754 0.977 0.981 0.969 0.886
egy 0.409 0.409 0.409 0.455 0.500 0.455 0.409 0.455
ell 0.931 0.937 0.985 0.676 0.948 0.975 0.984 0.842
eng 0.981 0.979 0.973 0.960 0.983 0.987 0.974 0.984
est 0.729 0.699 0.885 0.561 0.800 0.800 0.915 0.587
eus 0.916 0.927 0.963 0.952 0.916 0.938 0.982 0.905
fao 0.707 0.647 0.509 0.556 0.772 0.806 0.552 0.681
fas 0.756 0.810 0.567 0.577 0.837 0.919 0.565 0.655
fin 0.736 0.744 0.947 0.562 0.809 0.893 0.935 0.645
fra 0.994 0.963 0.963 0.984 0.993 0.989 0.976 0.928
frm 0.997 0.741 0.745 0.905 0.997 0.847 0.820 0.765
fro 0.782 0.701 0.686 0.709 0.822 0.725 0.694 0.679
gla 0.955 0.924 0.924 0.939 0.909 0.924 0.939 0.970
gle 0.750 0.821 0.750 0.786 0.679 0.750 0.750 0.714
glg 0.849 0.807 0.798 0.788 0.879 0.895 0.789 0.754
got 0.630 0.642 0.588 0.599 0.645 0.631 0.580 0.598
grc 0.824 0.719 0.683 0.623 0.887 0.758 0.711 0.707
guj 1.000 1.000 1.000 1.000 1.000 0.857 1.000 1.000
hbo 0.711 0.712 0.683 0.706 0.737 0.764 0.629 0.679
hbs 0.892 0.848 0.619 0.603 0.922 0.929 0.616 0.723
heb 0.853 0.829 0.609 0.642 0.876 0.868 0.585 0.667
hin 0.854 0.934 0.975 0.971 0.916 0.966 0.977 0.748
hit 0.620 0.640 0.520 0.600 0.560 0.480 0.540 0.540
hsb 0.683 0.677 0.624 0.629 0.667 0.694 0.624 0.591
hun 0.928 0.867 0.728 0.692 0.938 0.925 0.686 0.740
hye 0.883 0.898 0.631 0.623 0.929 0.946 0.662 0.671
hyw 0.813 0.781 0.583 0.608 0.894 0.859 0.551 0.629
isl 0.710 0.751 0.653 0.660 0.767 0.863 0.636 0.667
ita 0.925 0.910 0.915 0.670 0.952 0.965 0.915 0.791
Table 11: Multilingual benchmarking results on MultiBLIMP (ISO 639 language codes a*–i* in alphabetical order). “Ours” refers to PleIAs models pre-trained on Common Corpus. Within each model group, the best score is in bold, and the second-best is underlined.
Model Ours Gemma 3 XGLM BLOOM Ours Gemma 3 XGLM OLMo
350M 270M 564M 560M 1.2B 1B 1.7B 1B
kat 0.931 0.951 0.917 0.760 0.951 0.951 0.809 0.907
kaz 0.705 0.792 0.647 0.682 0.780 0.844 0.682 0.688
kir 0.930 0.843 0.914 0.919 0.935 0.924 0.843 0.941
kmr 0.710 0.662 0.588 0.579 0.761 0.748 0.577 0.588
koi 0.628 0.488 0.605 0.558 0.558 0.651 0.628 0.605
kpv 0.641 0.591 0.553 0.547 0.700 0.628 0.581 0.547
krl 0.650 0.612 0.688 0.573 0.612 0.642 0.704 0.573
kxh 0.483 0.433 0.475 0.333 0.458 0.450 0.442 0.483
lat 0.874 0.651 0.578 0.575 0.925 0.730 0.568 0.625
lav 0.791 0.747 0.616 0.604 0.844 0.862 0.611 0.623
lij 0.783 0.744 0.669 0.638 0.780 0.807 0.701 0.665
lit 0.928 0.848 0.745 0.740 0.947 0.932 0.736 0.779
mar 0.737 0.717 0.735 0.667 0.713 0.776 0.726 0.776
mdf 0.537 0.622 0.524 0.537 0.622 0.585 0.561 0.500
mkd 0.923 0.974 0.769 0.769 0.821 1.000 0.590 0.718
myv 0.608 0.614 0.565 0.560 0.636 0.619 0.532 0.547
nds 0.736 0.729 0.674 0.663 0.749 0.732 0.674 0.700
nhi 0.526 0.579 0.474 0.632 0.553 0.579 0.447 0.500
nld 0.924 0.912 0.620 0.627 0.954 0.963 0.663 0.829
olo 0.679 0.668 0.795 0.611 0.753 0.711 0.842 0.595
orv 0.733 0.721 0.690 0.636 0.757 0.744 0.707 0.667
ota 0.879 0.929 0.899 0.848 0.939 0.949 0.889 0.828
pcm 1.000 1.000 1.000 0.923 1.000 0.962 1.000 0.885
pol 0.892 0.849 0.624 0.634 0.930 0.931 0.628 0.725
por 0.948 0.933 0.939 0.955 0.965 0.972 0.920 0.872
quc 0.779 0.672 0.649 0.740 0.740 0.664 0.679 0.656
ron 0.868 0.874 0.638 0.608 0.903 0.928 0.640 0.793
rus 0.921 0.916 0.937 0.727 0.952 0.963 0.954 0.819
sah 0.688 0.771 0.736 0.792 0.708 0.681 0.701 0.764
san 0.657 0.666 0.612 0.609 0.670 0.678 0.618 0.620
slk 0.797 0.739 0.528 0.570 0.824 0.861 0.533 0.588
slv 0.882 0.796 0.618 0.636 0.903 0.854 0.622 0.711
sme 0.689 0.705 0.653 0.660 0.681 0.700 0.668 0.659
sms 0.833 0.802 0.779 0.757 0.821 0.779 0.764 0.768
spa 0.959 0.945 0.950 0.966 0.970 0.973 0.956 0.896
sqi 0.786 0.823 0.494 0.588 0.823 0.881 0.539 0.765
swe 0.995 0.995 0.970 0.950 0.995 1.000 0.985 0.990
tam 0.942 0.942 0.969 0.966 0.932 0.953 0.976 0.746
tpn 0.111 0.000 0.111 0.111 0.222 0.111 0.000 0.000
ttc 0.478 0.478 0.493 0.449 0.449 0.464 0.435 0.478
tur 0.766 0.804 0.829 0.700 0.814 0.908 0.857 0.716
uig 0.764 0.760 0.722 0.732 0.755 0.757 0.686 0.740
ukr 0.874 0.892 0.640 0.606 0.911 0.946 0.648 0.704
urb 0.538 0.462 0.462 0.231 0.538 0.462 0.462 0.462
urd 0.858 0.925 0.956 0.935 0.896 0.964 0.960 0.736
uzb 0.900 0.880 0.780 0.720 0.900 0.940 0.760 0.880
vep 0.572 0.588 0.588 0.481 0.631 0.626 0.567 0.524
wbp 0.250 0.000 0.167 0.250 0.250 0.083 0.250 0.250
wol 0.892 0.881 0.851 0.891 0.868 0.879 0.854 0.823
xcl 0.679 0.674 0.585 0.636 0.711 0.702 0.613 0.616
xnr 0.791 0.756 0.779 0.616 0.744 0.814 0.721 0.733
xpg 0.820 0.900 0.820 0.860 0.880 0.900 0.900 0.900
yrl 0.689 0.722 0.594 0.604 0.683 0.669 0.626 0.601
Table 12: Multilingual benchmarking results on MultiBLIMP (ISO 639 language codes k*–y* in alphabetical order). “Ours” refers to PleIAs models pre-trained on Common Corpus. Within each model group, the best score is in bold, and the second-best is underlined.
Model Ours Gemma 3 XGLM BLOOM Ours Gemma 3 XGLM OLMo
350M 270M 564M 560M 1.2B 1B 1.7B 1B
ar 0.475 0.492 0.500 0.521 0.477 0.572 0.525 0.473
ca 0.514 0.513 0.567 0.561 0.535 0.600 0.602 0.509
en 0.569 0.614 0.606 0.612 0.617 0.698 0.645 0.704
es 0.520 0.558 0.549 0.555 0.543 0.628 0.593 0.556
eu 0.516 0.524 0.531 0.538 0.514 0.531 0.561 0.503
gl 0.490 0.486 0.461 0.467 0.532 0.576 0.484 0.459
hi 0.509 0.541 0.520 0.549 0.515 0.598 0.557 0.493
id 0.500 0.544 0.542 0.555 0.518 0.631 0.583 0.498
my 0.492 0.507 0.515 0.475 0.498 0.518 0.537 0.475
ru 0.503 0.547 0.562 0.488 0.531 0.634 0.600 0.512
sw 0.500 0.507 0.531 0.500 0.510 0.551 0.563 0.494
te 0.542 0.562 0.559 0.557 0.553 0.592 0.581 0.532
zh 0.491 0.536 0.533 0.545 0.494 0.594 0.561 0.511
Table 13: Multilingual benchmarking results on XStoryCloze. “Ours” refers to PleIAs models pre-trained on Common Corpus. Languages are represented as two-letter codes in ISO 639.
Model Ours Gemma 3 XGLM BLOOM Ours Gemma 3 XGLM OLMo
350M 270M 564M 560M 1.2B 1B 1.7B 1B
es 0.566 0.590 0.604 0.620 0.614 0.678 0.664 0.524
et 0.518 0.498 0.554 0.488 0.500 0.536 0.568 0.480
eu 0.504 0.514 0.512 0.502 0.518 0.502 0.534 0.516
ht 0.522 0.504 0.548 0.500 0.524 0.518 0.556 0.534
id 0.534 0.578 0.574 0.596 0.558 0.690 0.646 0.544
it 0.542 0.524 0.536 0.502 0.562 0.648 0.536 0.488
qu 0.522 0.492 0.492 0.500 0.500 0.502 0.522 0.506
sw 0.528 0.546 0.530 0.516 0.538 0.544 0.562 0.510
ta 0.536 0.560 0.562 0.558 0.556 0.566 0.550 0.550
th 0.546 0.538 0.550 0.538 0.550 0.584 0.580 0.532
tr 0.530 0.566 0.544 0.528 0.542 0.606 0.536 0.530
vi 0.550 0.598 0.584 0.602 0.526 0.694 0.630 0.494
zh 0.536 0.564 0.554 0.588 0.544 0.640 0.584 0.522
Table 14: Multilingual benchmarking results on XCopa. “Ours” refers to PleIAs models pre-trained on Common Corpus. Languages are represented as two-letter codes in ISO 639.